### v1

# Deep Learning for 3D Human Pose Estimation and Mesh Recovery: A Survey

Abstract
3D human pose estimation and mesh recovery have attracted widespread research interest in many areas, such as computer vision, autonomous driving, and robotics. Deep learning on 3D human pose estimation and mesh recovery has recently thrived, with numerous methods proposed to address different problems in this area. In this paper, to stimulate future research, we present a comprehensive review of recent progress over the past five years in deep learning methods for this area by delving into over 200 references. To the best of our knowledge, this survey is arguably the first to comprehensively cover deep learning methods for 3D human pose estimation, including both single-person and multi-person approaches, as well as human mesh recovery, encompassing methods based on explicit models and implicit representations. We also present comparative results on several publicly available datasets, together with insightful observations and inspiring future research directions. A regularly updated project page can be found at https://github.com/liuyangme/SOTA-3DHPE-HMR.
keywords:
Human pose estimation, 3D human pose, human mesh recovery, human reconstruction, deep learning, literature survey
journal: Pattern Recognition
1 Introduction
1.1 Motivation
Humans are the foremost actors in various social activities, and it is imperative to equip AI with an understanding of humans for contributing to society. Consequently, there are many human-centric tasks positioning the epicenter of research. Among them, 3D Human Pose Estimation (HPE) and Human Mesh Recovery (HMR) represent crucial tasks in the field of computer vision to interpret the status and behavior of humans in complex real-world environments. 3D human pose estimation can accurately predict human body keypoint coordinates in three-dimensional space. This approach provides more comprehensive and accurate spatial information when compared to its 2D counterpart, thereby facilitating a better understanding of complex human behaviors in higher-level computer vision applications duan2022revisiting ; zhang2022voxeltrack ; zhu2022multilevel . Human mesh recovery reconstructs a three-dimensional digital model of the body, which captures details such as shape you2023co ; tripathi20233d , gestures fan2023hold , clothing dai2023cloth2body ; tang2023high , and facial expressions feng2023learning , thus offering direct insights into human interactions with the physical world. 3D pose estimation and mesh recovery have a broad range of applications, such as security and surveillance wang2021horeid , human-computer interaction liu2022arhpe , autonomous driving zheng2022multi ; wang2023learning , and virtual reality weng2019photo . With advanced deep learning technology, 3D human pose estimation and mesh recovery With advanced deep learning technology, 3D human pose estimation and mesh recovery have garnered increasing attention in recent years. 3D pose estimation has evolved from concentrating on single individuals to encompassing multiple persons with more varied data inputs. In human mesh recovery, advancements have been made in terms of data inputs and in capturing more intricate details. The augmentation in explicit model parameters allows for a more nuanced representation of the human body, and the expansion in parameter types facilitates the portrayal of finer surfaces. With the progress of implicit rendering and its incorporation into human mesh recovery, more flexible body representations are achieved. However, both 3D pose estimation and mesh recovery face significant challenges, such as multi-person scenarios, self-occlusion issues, and the detailed reconstruction of bodies. Thus, a systematic and comprehensive review of recent advancements in 3D human pose estimation and mesh recovery is essential.
1.2 Scope of this survey
Over the past five years, numerous reviews have been on 3D human pose estimation and mesh recovery. The reviews liu2022recent ; zheng2023deep primarily concentrate on 2D and 3D pose estimation, yet they devote less attention to HMR-related literature. The survey tian2023recovering provides a thorough review exclusively dedicated to mesh recovery, focusing on methods based on explicit models. The survey chen2021towards is distinctly centered on mesh recovery using implicit rendering techniques; however, it falls short in offering an exhaustive overview of the latest implicit rendering methods and systems for 3D pose estimation and mesh recovery. These reviews only marginally explore 3D pose estimation and mesh recovery technologies but scarcely delve into their applications in other computer vision tasks and future challenges.
This review primarily concentrates on deep learning approaches to 3D human pose estimation and human mesh recovery. 3D pose estimation, both in single-person and multi-person scenarios, is considered. In human mesh recovery, it methodically reviews techniques grounded in both explicit and implicit models. As illustrated in Fig.1, our survey comprehensively includes the most recent state-of-the-art publications (2019-2023) from mainstream computer vision conferences and journals.
Figure 1: Recent research of deep learning for 3D HPE and HMR.
The main contributions of this work compared to the existing literature can be summarized as follows: (1) To the best of our knowledge, this survey is arguably the first to comprehensively cover deep learning methods for 3D human pose estimation, including both single-person and multi-person approaches, as well as human mesh recovery, encompassing methods based on explicit models and implicit representations. (2) Unlike existing reviews, we have not overlooked the role of implicit representations in the methods, particularly with the recent rapid advances in implicit rendering. This approach can produce detailed outputs, including clothed human figures with expressions, movements, and other intricacies essential for achieving photorealism. (3) This paper comprehensively reviews the most recent developments in deep learning for 3D pose estimation and mesh recovery, providing readers with a detailed overview of cutting-edge methodologies. Additionally, it explores how these advancements contribute to various other computer vision tasks and delves into the challenges within this domain. The structure of this paper is as follows: Section 2 introduces the sensor type and representation for the human body that have been widely used. Section 3 presents the overview of deep learning for 3D human pose estimation and mesh recovery. Section 4 surveys existing single person and multi-person 3D pose estimation methods. Section 5 reviews the methods for human mesh recovery, including template-based and template-free methods. Section 6 introduces the evaluation metrics and datasets for the respective tasks. Moreover, Section 7 discusses the applications, along with their impact on other computer vision tasks. Finally, Section 8 concludes the paper. Additionally, we host a regularly updated project page, which can be accessed at: https://github.com/liuyangme/SOTA-3DHPE-HMR.
2 Background
2.1 Sensors used for 3D HPE and HMR
There are a variety of sensors that can be used for 3D human pose estimation and mesh recovery, which are mainly categorized as active sensors and passive sensors.
2.1.1 Active sensors
Active sensors operate by emitting a set of signals and measuring by detecting their reflections, such as Motion Capture (MoCap) systems with reflective markers einfalt2023uplift , tactile sensing luo2021intelligent , Time of Flight (ToF) cameras ruget2022pixels2pose ; pandey2019volumetric , and Radio Frequency (RF) technologies ren2022gopose ; li2022unsupervised . MoCap and tactile devices are only suitable for cooperative targets. Active cameras generally cannot be used outdoors, and using multiple active devices simultaneously may lead to mutual interference. Thus, these devices have limited application scenarios.
2.1.2 Passive sensors
Passive sensors do not actively emit any signals during the measurement; instead, they rely on signals from the objects or natural sources, including Inertial Measurement Units (IMUs) ponton2023sparseposer ; huang2020deepfuse and image sensors zou2022human ; xu2020eventcap . Among these, RGB image sensors are particularly notable for being simple, user-friendly, adaptable to various environments, and capable of capturing high-resolution color images. Additionally, multiple RGB image sensors can be combined to form multi-view systems jiang2023probabilistic ; shuai2022adaptive ; huang2021dynamic . In this survey, considering the widespread applicability of RGB image sensors and the length limitations of the article, we focus on 3D human pose estimation and mesh recovery using RGB image sensors.
2.2 Representation for human body
3D pose estimators and mesh reconstructors can generate corresponding outcomes from the sensor input described previously. The 3D human pose estimation output comprises 3D coordinates detailing positions and joint orientations of the human body, representing the spatial location of each joint (e.g., head, neck, shoulders, elbows, knees), and the skeleton maps the interconnections between joints. Typically, these coordinates are expressed in a global coordinate system or relative to the camera’s coordinate system. A keypoints tree structure is commonly employed to illustrate the human pose, where the tree nodes represent the joints and the edges denote the connections between joints. However, the skeleton-based human pose representation method does not provide detailed information about the body’s surface. The output of human mesh recovery is generally a 3D body model, which encompasses a comprehensive depiction of the body’s shape and surface details and offers a more enriched representation. Statistical-based models are widely used in human mesh representations, such as SCAPE anguelov2005scape and SMPL loper2015smpl . SMPL is a learnable skin-vertex model representing the human body as a 3D mesh with a topological structure. The pose and shape of the body are described by pose parameters $\theta$ and shape parameters $\beta$. The pose parameters control the joint angles and global posture, and the shape parameters determine the body’s shape. There are several models based on SMPL to expand the representational capabilities, such as the MANO model (SMPL+H) romero2022embodied with hand representation and the FLAME model li2017learning with facial representation. SMPL-X pavlakos2019expressive is a comprehensive model that simultaneously captures the human body, face, and hands by incorporating the FLAME head model li2017learning and the MANO hand model romero2022embodied . H4D jiang2022h4d can represent the dynamic human shape and pose by building upon the prior knowledge from the SMPL model and extending its capabilities by incorporating a temporal dimension. In addition to the explicit model representations mentioned above, recent years have seen the development of methods based on implicit models, which offer a more flexible representation of the human body through non-parametric models. For example, to address the limitations of previous representation models, Zheng et al. zheng2021pamir designed the Parametric Model-Conditioned Implicit Representation (PaMIR), which utilizes the 2D image feature map and corresponding SMPL feature volume to generate an implicit surface representation.
3 Overview of Deep Learning for 3D HPE and HMR
Figure 2: A basic framework of deep learning for 3D HPE and HMR.
This section provides an overview of the deep learning framework for 3D human pose estimation and mesh recovery. As shown in Fig.2, there are four components in deep learning-based systems. First, data are collected from various sensors, including RGB images, depth images, RF signals, IMUs, and so on. Second, the input data are processed through a deep learning model, which typically consists of an encoder and a decoder. The encoder extracts representational features from the input data, such as those obtained using architectures like ResNet he2016deep or HRNet wang2020deep . The decoder, which could be based on frameworks like MLP (Multi-Layer Perceptron) tolstikhin2021mlp or Transformer dosovitskiy2020image , then outputs the body’s 3D pose and reconstruction model. This structure enables the model to efficiently process and translate complex input data into detailed and accurate 3D representations of human pose and mesh. Thirdly, various learning methods can be selected during the model learning process. In addition to fully supervised learning, to alleviate data dependency approaches such as weakly supervised learning yang2023camerapose ; zanfir2020weakly , unsupervised learning chai2023global ; yu2021skeleton2mesh , and few-shot learning mu2023actorsnerf ; benzine2020pandanet are employed. To reduce the model size, techniques such as knowledge distillation yang2023effective ; tripathi2020posenet3d , model pruning liu2019effective , and parameter quantization choi2021mobilehumanpose can be applied. Furthermore, methodologies such as meta-learning cho2021camera and reinforcement learning gong2022posetriplet can also be incorporated, allowing the model to adapt to different scenarios and data constraints. Finally, the deep learning model outputs the results of 3D human pose estimation and mesh recovery. These results can be represented in various forms, including keypoints hassan2023regular ; zhao2023poseformerv2 , mesh loper2015smpl ; romero2022embodied ; cai2023smpler , and voxels li2020monocular ; onizuka2020tetratsdf ; tang2023high . These representations contribute to a comprehensive understanding of the human body in three dimensions. In the following sections, we will delve into the specifics of 3D human pose estimation and mesh recovery, categorizing and elaborating on each aspect.
4 3D Human Pose Estimation
3D human pose estimation can provide a more accurate pose by predicting the depth information of body keypoints, but it is much more challenging than 2D pose estimation. 3D pose estimation can be classified into single-person and multi-person estimation, according to the number of targets. Single-person 3D pose estimation has seen rapid progress due to the fast development of deep learning technology, but it still faces many challenges, such as efficiency and the invisibility of certain body parts. Multi-person estimation in crowded scenes is even more challenging because of the interaction and occlusion between bodies and scenes. In this section, we will detail the research progress in this field, categorizing it into two types. Table 1 summarizes all the representative methods.
Table 1: Overview of 3D human pose estimation.
Motivations Methods
Single Person in Images Solving Depth Ambiguity $\bullet$ Optical-aware: VI-HC wei2019view , Ray3D zhan2022ray3d
$\bullet$ Appropriate feature representation: HEMlets zhou2021hemlets
Solving Body Structure Understanding $\bullet$ Joint aware: JRAN zheng2020joint
$\bullet$ Limb aware: Wu et al. wu2021limb , Deep grammar network xu2021monocular
$\bullet$ Orientation keypoints: Fisch et al. fisch2021orientation
$\bullet$ Graph-based: Liu et al. liu2019feature , LCN ci2020locally , Modulated-GCN zou2021modulated , Skeletal-GNN zeng2021learning , HopFIR zhai2023hopfir , RS-Net hassan2023regular
Solving Occlusion Problems $\bullet$ Learnable-triangulation iskakov2019learnable
$\bullet$ RPSM qiu2019cross
$\bullet$ Lightweight multi-view remelli2020lightweight
$\bullet$ AdaFuse zhang2021adafuse
$\bullet$ Bartol et al. bartol2022generalizable
$\bullet$ 3D pose consensus luvizon2022consensus
$\bullet$ Probabilistic triangulation module jiang2023probabilistic
Solving Data Lacking $\bullet$ Unsupervised learning: Kudo et al. kudo2018unsupervised , Chen et al. chen2019unsupervised , ElePose wandt2022elepose
$\bullet$ Self-supervised learning: EpipolarPose kocabas2019self , Wang et al. wang20193d , MRP-Net kundu2022uncertainty , PoseTriplet gong2022posetriplet
$\bullet$ Weakly-supervised learning: Hua et al. hua2022weakly , CameraPose yang2023camerapose
$\bullet$ Transfer learning: Adaptpose gholami2022adaptpose
in Videos Solving Single-frame Limitation $\bullet$ VideoPose3D pavllo20193d
$\bullet$ PoseFormer zheng20213d
$\bullet$ UniPose+ artacho2021uniposeplus
$\bullet$ MHFormer li2022mhformer
$\bullet$ MixSTE zhang2022mixste
$\bullet$ Honari et al. honari2022temporal
$\bullet$ HSTFormer qian2023hstformer
$\bullet$ STCFormer tang20233d
Solving Real-time Problems $\bullet$ Temporally sparse sampling: Einfalt et al. einfalt2023uplift
$\bullet$ Spatio-temporal sparse sampling: MixSynthFormer sun2023mixsynthformer
Solving Body Structure Understanding $\bullet$ Motion loss: Wang et al. wang2020motion
$\bullet$ Human-joint affinity: DG-Net zhang2021learning
$\bullet$ Anatomy-aware: Chen et al. chen2021anatomy
$\bullet$ Part aware attention: Xue et al. xue2022boosting
Solving Occlusion Problems $\bullet$ Optical-flow consistency constraint: Cheng et al. cheng2019occlusion
$\bullet$ Multi-view: MTF-Transformer shuai2022adaptive
Solving Data Lacking $\bullet$ Unsupervised learning: Yu et al. yu2021towards
$\bullet$ Weakly-supervised learning: Chen et al. chen2019weakly
$\bullet$ Semi-supervised learning: MCSS mitra2020multiview
$\bullet$ Self-supervised learning: Kundu et al. kundu2020self , P-STMO shan2022p
$\bullet$ Meta-learning: Cho et al. cho2021camera
$\bullet$ Data augmentation: PoseAug gong2021poseaug , Zhang et al. zhang2023learning
Multi-person Top-down Solving Real-time Problems $\bullet$ Multi-view: Chen et al. chen2020cross
$\bullet$ Whole body: AlphaPose fang2022alphapose
Solving Representation Limitation $\bullet$ VoxelTrack zhang2022voxeltrack
Solving Occlusion Problems $\bullet$ Wu et al. wu2021graph
Solving Data Lacking $\bullet$ Single-shot: PandaNet benzine2020pandanet
$\bullet$ Optical-aware: Moon et al. moon2019camera
Bottom-up Solving Real-time Problems $\bullet$ Fabbri et al. fabbri2020compressed
Solving Supervisory Limitation $\bullet$ HMOR wang2020hmor
Solving Data Lacking $\bullet$ Single-shot: SMAP zhen2020smap , Benzine et al. benzine2021single
Solving Occlusion Problems $\bullet$ Mehta et al. mehta2018single
$\bullet$ LCR-Net++ rogez2019lcr
Others Single Stage $\bullet$ Jin et al. jin2022single
Top-down + Bottom-up $\bullet$ Cheng et al. cheng2022dual
4.1 Single person 3D pose estimation
As illustrated in Fig.3, single person 3D pose estimation is mainly classified into the direct estimation method and the 2D to 3D lifting method. The direct method estimates the 3D human pose directly from the input using a predictor, and the 2D to 3D lifting method, which estimates the 3D pose from the results of 2D estimation, involves first detecting the coordinates of human keypoints in 2D space, and then lifting these 2D keypoints onto 3D space coordinates.
(a) The direct estimation method
(b) The 2D to 3D lifting method
Figure 3: Typical single person 3D human pose estimation. (a) The direct estimation method; (b) The 2D to 3D lifting method.
4.1.1 Single person 3D pose estimation in images
Solving Depth Ambiguity. As shown in Fig.4(a), different 3D pose coordinates projecting to 2D images may give the same results, leading to an ill-posed problem. This problem can be solved by using the propagation properties of light and camera imaging principle. To address the ill-posed problem, Wei et al. wei2019view introduced a view-invariant framework to moderate the effects of viewpoint diversity. A View-Invariant Hierarchical Correction (VI-HC) network predicts the 3D pose refinement with view-consistent constraints in the proposed framework. Additionally, a view-invariant discriminative network actualizes high-level constraints after the base network generates an initial estimation. Ray-based 3D (Ray3D) absolute estimation method zhan2022ray3d can convert the input from pixel space to 3D normalized rays. Another helpful approach for this problem is learning a more appropriate feature representation. Part-centric HEatMap triplets (HEMlets) framework zhou2021hemlets utilizes three joint-heatmaps to represent the end-joints relative depth information, which bridges the gap between the 2D location and the 3D human pose.
(a) Depth ambiguity
(b) Graph-based representation
(c) Transfer learning
Figure 4: (a) Depth ambiguity; (b) Graph-based representation for human body; (c) Transfer learning.
Solving Body Structure Understanding. Unlike other computer vision tasks, the body’s unique structures can provide constraints or prior information to improve pose estimation performance. In the Joint Relationship Aware Network (JRAN) zheng2020joint , a dual attention module was designed to generate both whole and local feature attention block weights. The limb poses aware network wu2021limb leverages kinematic constraint and trajectory information to prevent errors from accumulating along the human body structure. Pose grammar in xu2021monocular learns a 2D-3D mapping function, enabling the input 2D pose to be transformed into 3D space, and integrates three aspects of human structure (kinematics, symmetry, motor coordination) into a Bi-directional Recurrent Neural Network (RNN). The orientation keypoints-based method fisch2021orientation uses virtual markers to generate sufficient information for accurately inferring rotations through simple post-processing. The Graph Neural Network (GNN) is a convolutional network defined on the graph data structure. It is difficult for typical neural networks to handle graph-structured data, such as body structure information. However, this challenge can be more easily addressed if graph-based learning methods are used, as shown in Fig.4(b). Liu et al. liu2019feature proposed a feature boosting network in which features are learned by the convolutional layers and are boosted with Graphical ConvLSTM to perceive the graphical long short-term dependency among different body parts. The Locally Connected Network (LCN) ci2020locally leverages the allocation of dedicated filters rather than sharing them for different joints. Modulated Graph Convolutional Network (Modulated-GCN) zou2021modulated includes weight and affinity modulation to learn modulation vectors for different body joints and to adjust the graph structure, respectively. Zeng et al. zeng2021learning designed a skeletal GNN learning framework to address the depth ambiguity and self-occlusion problems in 3D human pose estimation. In this framework, the proposed hop-aware hierarchical channel-squeezing fusion layer is designed to extract relevant information from neighboring nodes. Higher-order Regular Splitting graph Network (RS-Net) hassan2023regular captures long-range dependencies between body joints using multi-hop neighborhoods. It learns distinct modulation vectors for different body joints and adds modulation matrices to the corresponding skeletal adjacency matrices. Zhai et al. zhai2023hopfir employed intra-group joint refinement utilizing attention mechanisms to discover potential joint synergies to explore the potential synergies between joints.
Solving Occlusion Problems. Partial occlusion of the human body, including self-occlusion and other occlusions (such as object occlusion and multi-person occlusion), is typical in various scenes. Occlusion may interfere with pose estimation, potentially resulting in the prediction of an error pose. Solving the occlusion problem through a multi-view approach is effective, where an occluded pose not visible in one view is likely to be visible in other views in a multi-view system. Thus, a multi-view based approach can provide more reliable results through cross-view frame inference. In practice, Iskakov et al. iskakov2019learnable combined algebraic and volumetric triangulation for 3D pose estimation from multi-view 2D images. The former method is based on a differentiable algebraic triangulation with confidence weights, while the latter method utilizes volumetric aggregation from intermediate 2D backbone feature maps. Multi-view geometric priors have been incorporated in qiu2019cross , which combines two steps: first, predicting the 2D poses in multi-view RGB images through cross-view fusion, and second, utilizing the proposed Recursive Pictorial Structure Model (RPSM) to recover the 3D poses from the previously predicted multi-view 2D poses. To improve the real-time performance of multi-view 3D pose estimation, Remelli et al. remelli2020lightweight designed a lightweight framework with a differentiable Direct Linear Transform (DLT) layer. Adaptive multi-view fusion zhang2021adafuse enhances the features in occluded views by determining the correspondence between points in occluded and visible views. To address the problem of generalizability for multi-view estimation, Bartol et al. bartol2022generalizable proposed a stochastic learning framework for pose triangulation. Luvizon et al. luvizon2022consensus effectively integrated 2D annotated data and 3D poses to design a consensus-aware method, which optimizes multi-view poses from uncalibrated images by different coherent estimations up to a scale factor from the intrinsic parameters. Probabilistic triangulation module jiang2023probabilistic embedded in a 3D pose estimation framework can extend multi-view methods to uncalibrated scenes. It models camera poses using probability distributions and iteratively updates them from 2D features, replacing the traditional update through camera poses and eliminating the dependency on camera calibration.
Solving Data Lacking. In recent years, the development of diverse and efficient feature representations, along with end-to-end training modes, has significantly enhanced the accuracy of deep learning models in 3D human pose estimation. However, a significant limitation of these models is their reliance on fully supervised training, necessitating vast amounts of expensive and labor-intensive labeled 3D data, predominantly from indoor scenes. Addressing this challenge requires exploring alternative training strategies beyond fully supervised learning, such as unsupervised, semi-supervised, and few-shot learning approaches.
Unsupervised learning typically focuses on learning features rather than specific tasks from unlabeled training data, and it finds relationships between samples by mining the intrinsic features of the data. The first work kudo2018unsupervised can predict 3D human pose without any 3D dataset by adversarial learning, which is based on the generative adversarial networks via unsupervised training. The approach chen2019unsupervised utilizes geometric self-supervision and randomly reprojects 2D pose camera viewpoints from the recovered 3D skeleton during training, thereby forming a self-consistency loss in the lift-reproject-lift process. Elepose wandt2022elepose utilizes random projections, estimating likelihood using normalizing flows on 2D poses in a linear subspace. Furthermore, Elepose also learns the distribution of camera angles to reduce the dependence on camera rotation priors in the training dataset.
Self-supervised learning is a branch of unsupervised learning where the model can learn by itself from unlabeled training data and acquire the representation model on unlabeled data through pretext tasks. During on-the-ground execution, EpipolarPose kocabas2019self trains the 3D pose estimator without any 3D ground-truth data or camera extrinsic parameters, predicting 2D poses from multi-view images and obtaining a 3D pose and camera geometry via epipolar geometry. Wang et al. wang20193d designed a simple yet effective self-supervised correction mechanism for 3D human pose estimation by learning all intrinsic structures of the human pose, which is divided into two parts: the 2D-to-3D pose transformation and 3D-to-2D pose projection. MRP-Net kundu2022uncertainty reformulated the self-supervised 3D human pose estimation as an unsupervised domain adaptation problem, which includes model-free joint localization and model-based parametric regression. To reduce dependence on the consistency loss that guides learning, unlike previous works, the method gong2022posetriplet generates 2D-3D pose pairs for augmenting supervision via the proposed self-enhancing dual-loop learning framework.
A weakly-supervised model, characterized by its reliance on only weak labels, often faces the challenge of performing complex tasks with limited or imprecise guidance. Despite the lack of detailed annotations typically required in fully supervised scenarios, these models derive meaningful insights, utilizing vague or less informative labels to infer intricate patterns and relationships within the data. This ability to operate with minimal supervision makes them particularly useful when acquiring comprehensive labeled data is impractical or costly. In practical applications, Hua et al. hua2022weakly proposed a weakly-supervised method that initially lifts the 2D keypoints into coarse 3D poses across two views using triangulation and then refines the 3D pose by employing spatial configurations and cross-view correlations. Yang et al. yang2023camerapose designed a weakly-supervised framework that utilizes projection relationships to estimate 3D poses solely from 2D pose annotations under the condition of known camera parameters. Transfer learning enables the transfer of model parameters from the source domain to the target domain, allowing us to share learned model parameters with a new model, as shown in Fig.4(c). This approach speeds up and optimizes efficiency, eliminating the need to learn from scratch, as is common with most networks. For example, Adaptpose gholami2022adaptpose is an end-to-end cross-dataset adaptation 3D human pose estimation predictor that is implemented using transfer learning for limited data applications.
4.1.2 Single person 3D pose estimation in videos
With hardware development, image data are often acquired and processed in videos. The video frame data can provide more continuous information than a single-frame image, predicting human pose estimation in sequence through the spatio-temporal domain. In addition, optical flow and scene flow information can be extracted from videos to predict 3D human pose from data of multimodal tang2023ftcm ; zhao2023poseformerv2 . Solving Single-frame Limitation. Using continuous image frames from video can capture dynamic changes and temporal information. By analyzing this information, deep learning models can extract motion features and spatio-temporal relationships, thereby effectively improving the effectiveness of 3D human pose estimation. Pavllo et al. pavllo20193d proposed a 3D human pose estimation method for video that extracts temporal cues with dilated convolutions over 2D keypoint trajectories, and then they designed a semi-supervised method with back-projection to improve performance. PoseFormer zheng20213d as a spatial-temporal transformer-based framework predicts 3D body pose by analyzing human joint relations within each frame and their temporal correlations across frames. Based on previous work artacho2020unipose , unipose+ artacho2021uniposeplus leverages multi-scale feature representations to enhance the feature extractors in the framework’s backbone. This framework utilizes contextual information across scales and joint localization with Gaussian heatmap modulation to improve decoder estimation accuracy. Furthermore, Multi-Hypothesis transFormer (MHFormer) li2022mhformer learns spatio-temporal representations with multiple plausible pose hypotheses in a three-stage framework. In Mixed Spatio-Temporal Encoder (MixSTE) framework zhang2022mixste , the temporal transformer block learns the temporal motion of each joint and their inter-joint spatial correlation. Honari et al. honari2022temporal proposed an unsupervised feature extraction method using Contrastive Self-Supervised (CSS) learning to extract temporal information from videos. Extending this line of research, Hierarchical Spatial-Temporaltrans Formers (HSTFormer) qian2023hstformer utilizes spatial-temporal correlations of joints at different levels simultaneously, marking a first in studying hierarchical transformer encoders with multi-level fusion. Recently, Spatio-Temporal Criss-cross (STC) attention block tang20233d decomposes correlation learning into space and time to reduce the computational cost of the joint-to-joint affinity matrix.
Solving Real-time Problems. However, while various methods improve 3D pose estimation performance, they also introduce new problems and challenges in video-based estimation. For example, the continuous video frames add a substantial computational cost, which poses a challenge to the processing efficiency of the system and makes real-time processing more difficult. To reduce the total computational complexity, Einfalt et al. einfalt2023uplift designed a transformer-based scheme that uplifts dense 3D poses from temporally sparse 2D pose sequences by temporal upsampling within Transformer blocks. Sun et al. sun2023mixsynthformer reduced the complexity of attention calculation in Transformers through spatio-temporal sparse sampling, enabling the estimation of 3D human poses in video sequences on computationally constrained platforms.
Solving Body Structure Understanding. At the same time, the continuous motion information provided by video has led researchers to combine videos with human kinematics. Wang et al. wang2020motion designed a motion loss that computes the difference between the motion patterns of the predicted and ground truth keypoint trajectories, along with motion encoding, a simple yet effective representation of keypoint motion for 3D human pose estimation in videos. Dynamical Graph Network (DG-Net) zhang2021learning dynamically determines the human-joint affinity and adaptively predicts human pose via spatial-temporal joint relations in videos. To address the shortcoming of directly regressing the 3D joint locations, Chen et al. chen2021anatomy proposed an anatomy-aware estimation framework. This human skeleton anatomy-based framework includes a bone direction prediction network and a bone length prediction network, and it effectively utilizes bone features to bridge the gap between 2D keypoints and 3D joints. Xue et al. xue2022boosting designed a part-aware temporal attention module capable of extracting each part’s temporal dependency separately.
Solving Occlusion Problems. To address occlusion problems, video-based 3D human pose estimation enables the use of a multi-view approach and leverages the continuity of inputs between video frames to predict occluded body parts. Cheng et al. cheng2019occlusion introduced an occlusion-aware model that utilizes estimated 2D confidence heatmaps of keypoints and an optical-flow consistency constraint to generate a more complete 3D pose in occlusion scenes. Additionally, Multi-view and Temporal Fusing Transformer (MTF-Transformer) shuai2022adaptive fuses multi-view sequences in uncalibrated scenes for 3D human pose estimation in videos. To reduce dependency on camera calibration, the framework infers the relationship between pairs of views with a relative attention mechanism.
Solving Data Lacking. Human actions are inherently continuous, making video-based pose estimation critical for enhanced understanding. Compared with 3D annotated still images, high-quality, annotated 3D videos are relatively scarce. Nevertheless, the Internet abounds with a vast repository of unlabeled video data, the utilization of which for learning purposes holds considerable importance. Thus, video-based 3D human pose estimation to alleviate data dependency is essential. In practice, Yu et al. yu2021towards divided the unsupervised 3D pose estimation process into two sub-tasks: a scale estimation module and a pose lifting module. The scale estimation module optimizes the 2D input pose, while the pose lifting module maps the optimized 2D pose to its 3D counterpart. In weakly-supervised methods, Chen et al. chen2019weakly proposed a weakly-supervised method, employing a view synthesis framework and a geometry-aware representation. In this framework, the method leverages 2D keypoints for supervision and learns a shared 3D representation between viewpoints by synthesizing the human pose from one viewpoint to another. Semi-supervised learning utilizes only partially labeled data, consisting of a large amount of unlabeled data and a small amount of labeled data. In this context, Mitra et al. mitra2020multiview present a multi-view consistent semi-supervised method, which can regress 3D human pose from unannotated, uncalibrated, but synchronized multi-view videos. In self-supervised methods, Kundu et al. kundu2020self employed prior knowledge of the body skeleton and disentangles the inherent factors of variation through part-guided human image synthesis. Similarly, Shan et al. shan2022p randomly mask the body joints in spatial and temporal domains to better capture spatial and temporal dependencies. Meta-learning (a.k.a. ”learning to learn”) is a process where algorithms optimize their learning strategy based on experience. It involves training a model on various learning tasks, enabling it to learn new tasks more efficiently using fewer data samples. Cho et al. cho2021camera present an optimization-based meta-learning algorithm for 3D human pose estimation. This algorithm can adapt to arbitrary camera distortion by generating synthetic distorted data from undistorted 2D keypoints during model training. Apart from changing the learning supervision methods, data augmentation is also a practical approach to increasing the amount of data and enhancing the model’s generalization performance. PoseAug gong2021poseaug adjusts geometric factors such as posture, body shape, and viewpoint for model learning under the premise of the discriminative module controlling the feasibility of augmented poses. Zhang et al. zhang2023learning generates more diverse and challenging poses online.
4.2 Multi-person 3D pose estimation
As depicted in Fig.5, multi-person 3D pose estimation can be divided into two main categories: the bottom-up method (estimation + association) and the top-down method (detection + estimation). With a two-stage pipeline, the top-down method first detects every person in the input images and then extracts each person’s keypoints in the previously detected bounding box. The bottom-up method initially detects all keypoints in one stage and subsequently associates them with their respective persons. For instance, Cheng et al. cheng2022dual integrated both top-down and bottom-up approaches in multi-person pose estimation, complementing each other’s shortcomings. The top-down method shows robustness against potential erroneous bounding boxes, while the bottom-up network is more robust in handling scale variations. Finally, the 3D poses estimated from both top-down and bottom-up networks are fed into an integrated network to obtain the final 3D pose. Some single-stage methods and approaches combine the two methods as mentioned above. Unlike mainstream two-stage solutions for multi-person 3D human pose estimation, Jin et al. jin2022single introduced a single-stage method. This method expresses the 2D pose in the image plane and the depth information of a 3D body instance via the proposed decoupled representation and predicts the scale information of instances by extracting 2D pose features and enabling depth regression.
Figure 5: Typical multi-person 3D pose estimation.
4.2.1 Top-down methods
In top-down methods, AlphaPose fang2022alphapose predicts whole-body multi-person 3D human pose including face, body, hand, and foot. In the proposed framework, the symmetric integral keypoint regression module achieves fast and fine localization, the parametric pose non-maximum suppression module helps to eliminate redundant human detections, and the pose aware identity embedding module enables joint pose estimation and tracking. To achieve better real-time performance in multi-person 3D pose estimation, Chen et al. chen2020cross proposed a multi-view temporal consistency method in real-time from videos, which matches the 2D inputs with 3D poses directly in three-dimensional space and achieves over 150 FPS (frames per second) on a 12-camera setup and 34 FPS on a 28-camera setup. Additionally, choosing appropriate representations can effectively improve multi-person pose estimation results. For instance, Zhang et al. zhang2022voxeltrack employed a voxel representation to predict multi-person 3D pose and tracking, where the proposed voxel representation can determine whether each voxel contains a particular body joint. In multi-person scenarios, occlusion issues may be more severe and complex than in single-person scenarios. To address this challenge, Wu et al. wu2021graph utilized graph neural networks to boost information passing efficiency for multi-person, multi-view 3D human pose estimation. The multi-view matching graph module associates coarse cross-view poses within the networks, and the center refinement graph module further refines the results. Single-shot learning trains the model using significantly less data than is typically required for supervised learning, aiming to improve effectiveness. PandaNet benzine2020pandanet as an anchor-based model introduces a pose-aware anchor selection strategy to discard ambiguous anchors. Moon et al. moon2019camera proposed a top-down, camera distance-aware method for multi-person 3D pose estimation without relying on ground-truth information.
4.2.2 Bottom-up methods
MubyNet zanfir2018deep is a typical bottom-up, multi-task method for multi-person 3D pose estimation, which allows for training all component parameters. In the research on the lightweight top-down framework, Fabbri et al. fabbri2020compressed utilized high-resolution volumetric heatmaps to improve the estimation performance and designed a volumetric heatmap autoencoder that can compress the size of the representation. Designing better supervisory methods could also be effective, Wang et al. wang2020hmor introduced a novel supervisory approach named Hierarchical Multi-person Ordinal Relations (HMOR) using a monocular camera and designed a comprehensive top-level model to learn these ordinal relations, enhancing the accuracy of human depth estimation through a coarse-to-fine architecture. In the top-down estimation on single-shot learning, Mehta et al. mehta2018single inferred whole body pose under strong partial occlusions through occlusion-robust pose-maps in their proposed single-shot framework. Zhen et al. zhen2020smap designed a single-shot, bottom-up framework that estimates the absolute positions of multiple people by leveraging depth-related cues across the entire image. After their previous work benzine2020pandanet , Benzine et al. benzine2021single proposed a single-shot 3D human pose estimation method that predicts multi-person 3D poses without the need for bounding boxes. This method extends the Stacked Hourglass Network newell2016stacked to handle multi-person situations. To address the issue of occlusion, LCR-Net++ rogez2019lcr integrates adjacent pose hypotheses to predict the multi-person 2D and 3D poses without approximating initial human localization when a person is partially occluded or truncated by the image boundaries.
4.3 Summary of 3D pose estimation
Research in single-person 3D pose estimation predominantly focuses on some of the aforementioned critical issues. While there are commendable studies in this field, the challenges are yet to be fully resolved, necessitating more comprehensive research. In contrast to image-based methods, utilizing video-based methods is unequivocally seen as the trajectory of future advancements. The methodology adopted in multi-person 3D pose estimation must be tailored to the specific context. The current prevalent techniques can be categorized into two main approaches, each with inherent limitations. Top-down methods depend heavily on human detection and are susceptible to detection inaccuracies, often leading to unreliable pose estimations in environments with multiple people. Conversely, bottom-up methods, which operate independently of human detection and thus are immune to errors, face challenges in accurately processing all individuals in a scene concurrently, particularly affecting the detection of smaller-scale figures. In line with other domains in computer vision, the one-stage, end-to-end methods represent this field’s future direction.
5 3D Human Mesh Recovery
Human mesh recovery can be divided into two categories based on their representation models: template-based (parametric) methods and template-free (non-parametric) methods, as shown in Fig.6. Template-based human mesh recovery reconstructs predefined models (such as SCAPE anguelov2005scape , SMPL loper2015smpl ) by estimating the model’s parameters. In contrast, template-free human mesh recovery predicts the 3D body directly from input data without reliance on predefined models. The parametric approaches can be more robust than non-parametric methods with the templates’ prior knowledge, but their flexibility and detail are inherently limited even when extended with more parameters like SMPL+H romero2022embodied and SMPL+X cai2023smpler . Human mesh recovery is a crucial technique for digitizing the human body, posing significant challenges in computer vision and computer graphics due to the complex nature of geometric textures and color variations. Moreover, human mesh recovery faces challenges similar to 3D pose estimation, including environmental interference, multi-person scenarios, and occlusion issues. In this section, we will introduce the dominant methods based on these categories and challenges, with a comprehensive summary provided in Table 2.
(a) Template based human mesh recovery
(b) Template-free human mesh recovery
Figure 6: Typical human mesh recovery. (a) Template based human mesh recovery method; (b) Template-free human mesh recovery method.
Table 2: Overview of human mesh recovery.
Main ideas Methods
Template-based Naked Multimodal Methods $\bullet$ Hybrid annotations: Rong et al. rong2019delving
$\bullet$ Optical flow: DTS-VIBE li2022deep
$\bullet$ Silhouettes: LASOR yang2022lasor
$\bullet$ Cropped image and bounding box: CLIFF li2022cliff
Utilizing Attention Mechanism $\bullet$ Part-driven attention: PARE kocabas2021pare
$\bullet$ Graph attention: Mesh Graphormer lin2021mesh
$\bullet$ Spatio-temporal attention: MPS-Net wei2022capturing , PSVT qiu2023psvt
$\bullet$ Efficient architecture: FastMETRO cho2022cross , Xue et al. xue20223d
$\bullet$ End-to-end structure: METRO lin2021end
Exploiting Temporal Information $\bullet$ Temporally encoding features: Kanazawa et al. kanazawa2019learning
$\bullet$ Self-attention temporal: VIBE kocabas2020vibe
$\bullet$ Temporally consistent: TCMR choi2021beyond
$\bullet$ Multi-level spatial-temporal attention: MAED wan2021encoder
$\bullet$ Temporally embedded live stream: TePose wang2022live
$\bullet$ Short-term and long-term temporal correlations: GLoT shen2023global
Multi-view Methods $\bullet$ Confidence-aware majority voting mechanism: Dong et al. dong2021shape
$\bullet$ Probabilistic-based multi-view: Sengupta et al. sengupta2021probabilistic
$\bullet$ Dynamic physics-geometry consistency: Huang et al. huang2021dynamic
$\bullet$ Cross-view fusion: Zhuo et al. zhuo2023towards
Boosting Efficiency $\bullet$ Sparse constrained formulation: SCOPE fan2021revitalizing
$\bullet$ Single-stage model: BMP zhang2021body
$\bullet$ Process heatmap inputs: HeatER zheng2022heater
$\bullet$ Removing redundant tokens: TORE dou2023tore
Developing Various Representations $\bullet$ Texture map: TexturePose pavlakos2019texturepose
$\bullet$ UV map: Zhang et al. zhang2020learning , DecoMR zeng20203d , Zhang et al. zhang2020object
$\bullet$ Heat map: Sun et al. sun2021monocular , 3DCrowdNet choi2022learning
$\bullet$ Uniform representation: DSTformer zhu2023motionbert
Utilizing Structural Information $\bullet$ Part-based: holopose guler2019holopose
$\bullet$ Skeleton disentangling: Sun et al. sun2019human
$\bullet$ Hybrid inverse kinematics: HybrIK li2021hybrik , NIKI li2023niki
$\bullet$ Uncertainty-aware: Lee et al. lee2021uncertainty
$\bullet$ Kinematic tree structure: Sengupta et al. sengupta2021hierarchical
$\bullet$ Kinematic chains: SGRE wang20233d
Choosing Appropriate Learning Strategies $\bullet$ Self-improving: SPIN kolotouros2019learning , ReFit wang2023refit , You et al. you2023co
$\bullet$ Novel losses: Zanfir et al. zanfir2020weakly , Jiang et al. jiang2020coherent
$\bullet$ Unsupervised learning: Madadi et al. madadi2021deep , Yu et al. yu2021skeleton2mesh
$\bullet$ Bilevel online adaptation: Guan et al. guan2022out
$\bullet$ Single-shot: Pose2UV huang2022pose2uv
$\bullet$ Contrastive learning: JOTR li2023jotr
$\bullet$ Domain adaptation: Nam et al. nam2023cyclic
Detailed With Clothes $\bullet$ Alldieck et al. alldieck2019learning
$\bullet$ Multi-Garment Network (MGN) bhatnagar2019multi
$\bullet$ Texture map: Tex2Shape alldieck2019tex2shape
$\bullet$ Layered garment representation: BCNet jiang2020bcnet
$\bullet$ Temporal span: H4D jiang2022h4d
With Hands $\bullet$ Linguistic priors: SGNify forte2023reconstructing
$\bullet$ Two-hands interaction: zhang2021interacting
$\bullet$ Hand-object interaction: chen2021joint
Whole Body $\bullet$ PROX hassan2019resolving
$\bullet$ ExPose choutas2020monocular
$\bullet$ FrankMocap rong2021frankmocap
$\bullet$ PIXIE feng2021collaborative
$\bullet$ Moon et al. moon2022accurate
$\bullet$ PyMAF zhang2023pymaf
$\bullet$ OSX lin2023one
$\bullet$ HybrIK-X li2023hybrik
Template-free Regression-based $\bullet$ FACSIMILE smith2019facsimile , PeeledHuman jinka2020peeledhuman , GTA zhang2023global , NSF xue2023nsf
Optimization-based Differentiable $\bullet$ DiffPhy gartner2022differentiable , AG3D dong2023ag3d
Implicit Representations $\bullet$ PIFu saito2019pifu , PIFuHD saito2020pifuhd
$\bullet$ Canonical space: ARCH huang2020arch , ARCH++ he2021arch++ , CAR liao2023high
$\bullet$ Geometric priors: GeoPIFu he2020geo
$\bullet$ Novel representations: Peng et al. peng2021neural , 3DNBF zhang20233d
Neural Radiance Fields $\bullet$ Volume deformation scheme gao2022mps
$\bullet$ ActorsNeRF mu2023actorsnerf
Diffusion Models $\bullet$ HMDiff foo2023distribution
Implicit + Explicit $\bullet$ HMD zhu2019detailed , IP-Net bhatnagar2020combining , PaMIR zheng2021pamir , Zhu et al. zhu2021detailed , ICON xiu2022icon , ECON xiu2023econ , DELTA feng2023learning , GETAvatar zhang2023getavatar
Diffusion + Explicit $\bullet$ DINAR svitov2023dinar
NeRF + Explicit $\bullet$ TransHuman pan2023transhuman
Gaussian Splatting + Explicit $\bullet$ Animatable 3D Gaussian liu2023animatable
5.1 Template-based human mesh recovery
5.1.1 Naked human body recovery
Multimodal Methods. Multimodality in human mesh recovery harnesses the potential by combining various modalities of data, such as RGB images, depth information, and optical flow. Integrating data in different modalities can significantly enhance the robustness and precision of mesh recovery. Rong et al. rong2019delving proposed a hybrid annotation method with a hybrid training strategy for human mesh recovery to reduce annotation costs, which effectively utilizes various types of heterogeneous annotations, including 3D and 2D annotations, body part segmentation, and dense correspondence. Deep Two-Stream Video Inference for Human Body Pose and Shape Estimation (DTS-VIBE) li2022deep method redefines the task as a multimodal problem, merging RGB data with optical flow to achieve a more reliable estimation and address temporal inconsistencies from RGB videos. LASOR yang2022lasor estimates 3D pose and shape by synthesizing occlusion-aware silhouettes and 2D keypoints data in scenes with inter-person occlusion. CLIFF li2022cliff utilizes cropped images and bounding box information from the pre-training phase as inputs to enhance the accuracy of global rotation estimation in the camera coordinate system.
Utilizing Attention Mechanism. Transformer vaswani2017attention as a self-attention model has demonstrated remarkable success in Natural Language Processing (NLP) devlin2018bert brown2020language . Following this, the Vision Transformer (ViT) dosovitskiy2020image has mirrored these successes in the field of computer vision tolstikhin2021mlp liu2021swin , and numerous Transformer-based methods are now being employed in mesh recovery. The attention mechanism serves to amplify the importance of certain parts in the neural network. To address the partial occlusion issues, Part Attention REgressor (PARE) kocabas2021pare leverages the relationships between body parts derived from segmentation masks, prompting the network to enhance predictions for occluded body parts. Mesh Graphormer lin2021mesh utilizes a GCNN-reinforced transformer for estimating 3D mesh vertices and body joints, while a GCNN is employed to infer interactions among neighboring vertices based on pre-existing mesh topology. The architecture effectively merges graph-based networking with the attention mechanism of transformers, enabling the modeling of local and global interactions. MPS-Net wei2022capturing employs motion continuity attention to capture temporal coherence. This approach then leverages a hierarchical attentive feature mechanism to combine features from temporally adjacent representations more effectively. FastMETRO cho2022cross leverages self-attention mechanisms for non-adjacent vertices based on the topology of the body’s triangular mesh to minimize memory overhead and accelerate inference speed. Xue et al. xue20223d proposed a learnable sampling module for human mesh recovery to reduce inherent depth ambiguity. This module aggregates global information by generating joint adaptive tokens, utilizing non-local information within the input image. METRO lin2021end is a mesh transformer for end-to-end human mesh recovery, in which the encoder captures interactions between vertices and joints, and the decoder outputs 3D joint coordinates and the mesh structure. Qiu et al. qiu2023psvt proposed an end-to-end Transformer-based method for multi-person human mesh recovery in videos. This approach utilizes a spatio-temporal encoder to extract global features, followed by a spatio-temporal pose and shape decoder to predict human pose and mesh.
Exploiting Temporal Information. With the advancement of video technology, video-based human mesh recovery methods that extract temporal information from adjacent frames have shown increased potential. To learn the 3D dynamics of the human body more accurately and effectively, Kanazawa et al. kanazawa2019learning proposed a framework that produces smooth 3D meshes from videos. This framework also predicts past and future 3D motions by temporally encoding image features. Video Inference for Body pose and shape Estimation (VIBE) kocabas2020vibe estimates kinematically plausible motion sequences through self-attention mechanism-based temporal network and adversarial training without requiring any ground-truth 3D labels. To diminish the dependency on static features in the current frame, addressing a limitation of previous temporal-based methods, Choi et al. choi2021beyond developed the Temporally Consistent Mesh Recovery (TCMR) system. The system utilizes temporal information to ensure consistency and effectively recovers smooth 3D human motion by incorporating data from past and future frames. Multi-level Attention Encoder-Decoder (MAED) network wan2021encoder captures relationships at multiple levels, including the spatial-temporal and human joint levels, through its multi-level attention mechanism. Wang et al. wang2022live introduced the Temporally embedded 3D human body Pose and shape estimation (TePose) method, specifically tailored for live stream videos. They designed a motion discriminator for adversarial training, utilizing datasets without any 3D labels, through a multi-scale spatio-temporal graph convolutional network. Additionally, they employed a sequential data loading strategy to accommodate the unique start-to-end data processing requirements of live streaming. To effectively balance the learning of short-term and long-term temporal correlations, Global-to-Local Transformer (GLoT) shen2023global structurally decouples the modeling of long-term and short-term correlations.
Multi-view Methods. Dong et al. dong2021shape designed a practical multi-view framework that combines 2D observations from multi-view images into a unified 3D representation for individual instances using a confidence-aware majority voting mechanism. Sengupta et al. sengupta2021probabilistic introduced a probabilistic-based multi-view method without constraints such as specific target poses, viewpoints, or background conditions across image sequences. Huang et al. huang2021dynamic proposed a dynamic physics-geometry consistency approach for multi-person multi-view mesh recovery. This method integrates motion priors, extrinsic camera parameters, and human mesh data to mitigate the impact of noisy human semantic data. Zhuo et al. zhuo2023towards proposed a cross-view fusion method that predicts foot posture by achieving a more refined 3D intermediate representation and alleviating inconsistencies across different views.
Boosting Efficiency. Maintaining excellent performance with a lightweight model and low computational cost is essential and challenging, especially in applications such as wearable devices, power-limited systems, and edge computing. SCOPE fan2021revitalizing efficiently computes the Gauss-Newton direction using 2D and 3D keypoints for human mesh recovery. The method capitalizes on inherent sparsity and employs a sparse constrained formulation, achieving real-time performance at over 30 FPS. To implement a single-stage multi-person human mesh recovery model, Body Meshes as Points (BMP) zhang2021body represents multiple people as points and associates each with a single body mesh to significantly improve efficiency and performance. HeatER zheng2022heater processes heatmap inputs directly to reduce memory and computational costs. Dou et al. dou2023tore designed an efficient transformer for human mesh recovery to reduce model complexity and computational cost by removing redundant tokens.
Developing Various Representations. Stable and effective feature representation is crucial for enhancing the capabilities of deep learning algorithms in human mesh recovery, as it enables the efficient extraction of meaningful patterns from complex input data. Based on the assumption that image texture remains constant across frames, TexturePose pavlakos2019texturepose leverages the appearance constancy of the body across different frames. It measures the texture map for each frame using a texture consistency loss. DenseRaC xu2019denserac generates a pixel-to-surface correspondence map to optimize the estimation of parameterized human pose and shape. Zhang et al. zhang2020learning established a connection between 2D pixels and 3D vertices by using dense correspondences of body parts, effectively addressing related issues. Their proposed DaNet model concentrates on learning the 2D-to-3D mapping, while the PartDrop strategy ensures that the model focuses more on complementary body parts and adjacent positional features. To address the lack of dense correspondence between image features and the 3D mesh, DecoMR zeng20203d recovers the human mesh by establishing a pixel-to-surface dense correspondence map in the UV space. Zhang et al. zhang2020object designed a two-branch network that utilizes a partial UV map to represent the human body when occluded by objects, effectively converting this map into an estimation of the 3D human shape. Sun et al. sun2021monocular proposed a body-center-guided representation method that predicts both the body center heatmap and the mesh parameter map. This approach describes the 3D body mesh at the pixel level, enabling one-stage multi-person 3D mesh regression. 3DCrowdNet choi2022learning employs a joint-based regressor to isolate target features from others for recovering multi-person meshes in crowded scenes, which utilizes a crowded scene-robust image feature heatmap instead of the full feature map within a bounding box. Dual-stream Spatio-temporal Transformer (DSTformer) zhu2023motionbert extracts long-range spatio-temporal relationships among skeletal joints to effectively capture unified human motion representations from large-scale and heterogeneous video data for human-centric tasks, such as human mesh recovery.
Utilizing Structural Information. The structural information of the body acts as unique prior knowledge in human mesh recovery, enhancing the understanding of body relations. Furthermore, introducing additional physical constraints, which describe the interrelationships between various body structures, can significantly improve the performance. Holopose guler2019holopose employs a part-based multi-task regression network for 2D, 3D, and dense pose to estimate 3D human surfaces. Sun et al. sun2019human introduced an end-to-end method for human mesh recovery from single images and monocular videos. This method employs skeleton disentangling to reduce the complexity of decoupling and incorporates temporal coherence, efficiently capturing both short and long-term temporal cues. Hybrid Inverse Kinematics (HybrIK) li2021hybrik calculates the swing rotation from 3D joints and employs a network to predict the twist rotation through the twist-and-swing decomposition. The method can tackle non-linearity challenges and misalignment between images and models in mesh estimation from 3D poses. In their later work, Li et al. li2023niki designed NIKI, a model capable of learning from both the forward and inverse processes using invertible networks. To address the non-linear mapping and drifting joint position issues, Lee et al. lee2021uncertainty introduced an uncertainty-aware method for human mesh recovery, leveraging information from 2D poses to address the inherent ambiguities in 2D. Sengupta et al. sengupta2021hierarchical presented a probabilistic approach to circumvent the ill-posed problem, which integrates the kinematic tree structure of the human body with a Gaussian distribution over SMPL parameters. This method then predicts the hierarchical matrix-fisher distribution of 3D joint rotation matrices. SGRE wang20233d estimates the global rotation matrix of joints directly to avoid error accumulation along the kinematic chains in human mesh recovery.
Choosing Appropriate Learning Strategies. SPIN kolotouros2019learning incorporates an initial estimate optimization routine into the training loop by the self-improving neural network, which can fit the body mesh estimate to 2D joints. Zanfir et al. zanfir2020weakly developed a method integrating kinematic latent normalizing flow representations and dynamical models with structured, differentiable, semantic body part alignment loss functions aimed at enhancing semi-supervised and self-supervised 3D human pose and shape estimation. Jiang et al. jiang2020coherent introduced two novel loss functions for multi-person mesh recovery from single images: a distance field-based collision loss penalizing interpenetration between constructed figures and a depth ordering-aware loss addressing occlusions and promoting accurate depth ordering of targets. Madadi et al. madadi2021deep presented an unsupervised denoising autoencoder network to recover invisible landmarks using sparse motion capture data effectively. To tackle the challenges of pose failure and shape ambiguity in the unsupervised human mesh recovery task, Yu et al. yu2021skeleton2mesh devised a strategy that decouples the task into unsupervised 3D pose estimation and leverages kinematic prior knowledge. Bilevel Online Adaptation (BOA) guan2021bilevel employs bilevel optimization to reconcile conflicts between 2D and temporal constraints in out-of-domain streaming videos human mesh recovery. In their later work, Dynamic Bilevel Online Adaptation (DBOA) guan2022out integrates temporal constraints to compensate for the absence of 3D annotations. Huang et al. huang2022pose2uv developed Pose2UV, a single-shot human mesh recovery method capable of extracting target features under occlusions using a deep UV prior. JOTRli2023jotr fuses 2D and 3D features and incorporates supervision for the 3D feature through a Transformer-based contrastive learning framework. ReFit wang2023refit reprojects keypoints and refines the human model via a feedback-update loop mechanism. You et al. you2023co introduced a co-evolution method for human mesh recovery that utilizes 3D pose as an intermediary. This method divides the process into two distinct stages: initially, it estimates the 3D human pose from video, and subsequently, it regresses mesh vertices based on the estimated 3D pose, combined with temporal image features. To bridge the gap between training and test data, CycleAdapt nam2023cyclic proposed a domain adaptation method including a mesh reconstruction network and a motion denoising network enabling more effective adaptation.
5.1.2 Detailed human body recovery
Model-based human mesh recovery has moderately satisfactory results but still lacks detailed body representations; expansions on the naked parametric model now allow for parameterized depictions of various body parts and details, including clothing jiang2020bcnet ; jiang2022h4d , hands romero2022embodied , face li2017learning , and the entire body pavlakos2019expressive , as discussed in Section 2.2.
With Clothes. Alldieck et al. alldieck2019learning estimated the parameters of the SMPL model, including clothing and hair, from 1 to 8 frames of a monocular video. Bhatnagar et al. bhatnagar2019multi developed the Multi-Garment Network (MGN) to reconstruct body shape and clothing layers on top of the SMPL model. Tex2Shape alldieck2019tex2shape converts the human body mesh regression problem into an image-to-image estimation task. Specifically, it predicts a partial texture map of the visible region and then reconstructs the body shape, adding details to visible and occluded parts. BCNet jiang2020bcnet features a layered garment representation atop the SMPL model and innovatively decouples the skinning weight of the garment from the body mesh. Jiang et al. jiang2022h4d introduced a method that employs a temporal span, SMPL parameters of shape and initial pose, and latent codes encoding motion and auxiliary information. This approach facilitates the recovery of detailed body shapes, including visible and occluded parts, by utilizing a texture map of the visible region.
With Hands. Forte et al. forte2023reconstructing proposed SGNify, a model that captures hand pose, facial expression, and body movement from sign language videos. It employs linguistic priors and constraints on 3D hand pose to effectively address the ambiguities in isolated signs. Additionally, the relationship between Two-Hands zhang2021interacting , and Hand-Object chen2021joint effectively reconstructs the hand’s details.
Whole Body. To address the inconsistency between 3D human mesh and 3D scenes, Hassan et al. hassan2019resolving introduced a human whole body and scenes recovery method named Proximal Relationships with Object eXclusion (PROX). EXpressive POse and Shape rEgression (ExPose) framework choutas2020monocular employs a body-driven attention mechanism and adopts a regression approach for holistic expressive body reconstruction to mitigate local optima issues in optimization-based methods. FrankMocap rong2021frankmocap operates by independently running 3D mesh recovery regression for face, hands, and body and subsequently combining the outputs through an integration module. PIXIE feng2021collaborative integrates independent estimates from the body, face, and hands using the shared shape space of SMPL-X across all body parts. Moon et al. moon2022accurate developed an end-to-end framework for whole-body human mesh recovery named Hand4Whole, which employs joint features for 3D joint rotations to enhance the accuracy of 3D hand predictions. Zhang et al. zhang2023pymaf enhanced the PyMAF framework zhang2021pymaf , developing PyMAF-X for the detailed reconstruction of full-body models. This advancement aims to resolve the misalignment issues in regression-based, one-stage human mesh recovery methods by employing a feature pyramid approach and refining the mesh-image alignment parameters. OSX lin2023one employs a simple yet effective component-aware transformer that includes a global body encoder and a local face/hand decoder instead of separate networks for each part. Li et al. li2023hybrik extended the HybrIK li2021hybrik framework and proposed HybrIK-X, a one-stage model based on a hybrid analytical-neural inverse kinematics framework to recover comprehensive whole-body meshes with details.
5.2 Template-free human body recovery
Template-free methods for human mesh recovery, such as neural network regression models, optimization models based on differentiable rendering, implicit representation models, Neural Radiance Fields (NeRF), and Gaussian Splatting, demonstrate enhanced flexibility over template-based approaches, enabling the depiction of richer details.
(a) NeRF
(b) 3D Guassian Splatting
Figure 7: (a) NeRF; (b) 3D Guassian Splatting.
Regression-based human mesh recovery bypasses the limitations and biases inherent in template-based models and directly outputs the template-free 3D models of the human body. This approach allows for a more dynamic and flexible generation of models, which can capture a wider variety of human body shapes and postures. FACSIMILE (FAX) utilizes an image-translation network to recover geometry at the original image resolution directly, bypassing the need for indirect output through representations. Jinka et al. jinka2020peeledhuman proposed a robust shape representation specifically for scenes with self-occlusions, where the method encodes the human body using peeled RGB and depth maps, significantly enhancing accuracy and efficiency during both training and inference. Neural Surface Fields (NSF) xue2023nsf models a continuous and flexible displacement field on the base surface for 3D clothed human mesh recovery from monocular depth. This approach adapts to base surfaces with varying resolutions and topologies without retraining during inference. Zhang et al. zhang2023global introduced the Global-correlated 3D-decoupling Transformer for clothed Avatar reconstruction (GTA), a model that employs an encoder to capture globally-correlated image features and a decoder to decouple tri-plane features using cross-attention. Optimization-based differentiable rendering integrates the rendering process into the optimization method by minimizing the rendering error. AG3D dong2023ag3d captures the body’s and loose clothing’s shape and deformation by adopting a holistic 3D generator and integrating geometric cues in the form of predicted 2D normal maps. Gärtner et al. gartner2022differentiable designed DiffPhy, a differentiable physics-based model that incorporates a physically plausible body representation with anatomical joint limits, significantly enhancing the performance and robustness of human body recovery.
Implicit representations do not directly represent an object’s geometric data, such as vertices or meshes, but rather define whether a point in space belongs to the object through a function. The advantage of implicit representations lies in their ability to compactly and flexibly represent complex shapes, including those with intricate topologies or discontinuities. Thus, they have achieved impressive outcomes in human body reconstruction. Saito et al. saito2019pifu proposed the Pixel-aligned Implicit Function (PIFu), an implicit representation that aligns pixels of 2D images with a 3D object, and designed an end-to-end deep learning framework for inferring both 3D surface and texture from a single image. PIFu was the first to apply implicit representation in human mesh recovery, enabling the reconstruction of geometric details of the human body in clothing. Subsequently, PIFuHD saito2020pifuhd extended the work of PIFu to 4K resolution images, further enhancing detail. Huang et al. huang2020arch designed the Animatable Reconstruction of Clothed Humans (ARCH) method by creating a semantic space and a semantic deformation field. He et al. he2021arch++ introduced ARCH++, a co-supervising framework with cross-space consistency, which jointly estimates occupancy in both the posed and canonical spaces. They transform the human mesh recovery problem with implicit representation from pose space to canonical space for processing. However, this approach’s limitation is that it heavily relies on accurate pose estimation, and the clothing expression based on skinning weights still lacks sufficient naturalness in detail. GeoPIFu he2020geo learns latent voxel features using a structure-aware 3D U-Net incorporating geometric priors. To tackle the ill-posed problem in representation learning when views are incredibly sparse, Peng et al. peng2021neural developed a novel human body representation. This representation assumes that the neural representations learned at different frames share latent codes anchored to a deformable mesh. Clothed Avatar Reconstruction (CAR) method liao2023high employs a learning-based implicit model to initially form the general shape in canonical space, followed by refining surface details through predicting non-rigid optimization deformation in the posed space. 3D-aware Neural Body Fitting (3DNBF) zhang20233d addresses the challenges of occlusion and 2D-3D ambiguity by a volumetric human representation using Gaussian ellipsoidal kernels. As illustrated in Fig.7(a), Neural Radiance Fields mildenhall2021nerf is based on the principle of implicit representation, using neural networks to learn the continuous volumetric density and color distribution of a scene, allowing for generating a high-quality 3D model from arbitrary viewpoints. Gao et al. gao2022mps utilized a specialized representation that combines a canonical Neural Radiance Field (NeRF) with a volume deformation scheme, enabling the recovery of novel views and poses for a person not seen during training. Mu et al. mu2023actorsnerf introduce ActorsNeRF, a NeRF-based human representation for human mesh recovery from a few monocular images by encoding the category-level prior through parameter sharing with a 2-level canonical space. Diffusion models are based on a series of diffusion processes, transforming original data by adding random noise and then gradually removing this noise to generate new data through a reverse process. Human Mesh Diffusion (HMDiff) foo2023distribution treats mesh recovery as a reverse diffusion process, incorporates input-specific distribution information into the diffusion process, and introduces prior knowledge to simplify the task.
Methods based on implicit representations can recover free-form geometric shapes, but they may have excessive dependence on accurate poses or generate unsatisfactory shapes for novel poses or clothing. To increase robustness in these scenarios, existing work employs explicit body models to constrain mesh reconstruction in implicit methods. Researchers desire an approach that combines implicit and explicit methods based on a coarse body mesh’s geometry to refine and produce detailed human models meticulously. Hierarchical Mesh Deformation (HMD) zhu2019detailed uses constraints from body joints, silhouettes, and per-pixel shading information to combine the robustness of a parametric model with the flexibility of free-form 3D deformation. Bhatnagar et al. bhatnagar2020combining introduced an Implicit Part Network (IP-Net) to predict the detailed human surface, including the 3D surface of the dressed person, the inner body surface, and the parametric body model. Parametric Model-Conditioned Implicit Representation (PaMIR) zheng2021pamir combines a parametric body model with a free-form deep implicit function to enhance generalization through regularization. Zhu et al. zhu2021detailed utilized a ’project-predict-deform’ strategy that refines the SMPL model generated by human recovery methods using supervision from joints, silhouettes, and shading information. ICON xiu2022icon generates a stable, coarse human mesh using the SMPL model, then renders the front and back body normals, which provide rich texture details and combine with the original image. Finally, the body’s normal clothed normal, and Signed Distance Function (SDF) information are fed into an implicit MLP to obtain the final result. Subsequently, Xiu et al. proposed ECON xiu2023econ , which integrates normal estimation, normal integration, and shape completion by using SMPL-X depth as a soft geometric constraint within the optimization equation of Normal Integration. It endeavors to maintain coherence with nearby surfaces during the integration of normals. Feng et al. feng2023learning presented Disentangled Avatars (DELTA), a model representing humans with hybrid explicit-implicit 3D representations. GETAvatar zhang2023getavatar directly produces explicit textured 3D human meshes. GETAvatar creates an articulated 3D human representation with explicit surface modeling, enriches it with realistic surface details derived from 2D normal maps of 3D scan data, and utilizes a rasterization-based renderer for surface rendering. Diffusion Inpainting of Neural Avatars (DINAR) svitov2023dinar combines neural textures with the SMPL-X body model and employs a latent diffusion model to recover textures of both seen and unseen regions, subsequently integrating these onto the base SMPL-X mesh. TransHuman pan2023transhuman employs transformers to project the SMPL model into a canonical space and associates each output token with a deformable radiance field. This field encodes the query point in the observation space and is further utilized to integrate fine-grained information from reference images. Gaussian Splatting kerbl20233d is a technique for processing and rendering point clouds, using a Gaussian function to create an influence range for each point, resulting in a smoother and more natural representation in two-dimensional images, as shown in Fig.7(b). It has achieved good results in SLAM (Simultaneous Localization and Mapping) yan2023gs , generative human modeling liu2023humangaussian , dynamic scene reconstruction wu20234d , and multimodal generation chen2023text . Reconstruction based on Gaussian Splatting requires an initial point cloud input, which poses a challenge for human mesh recovery from image inputs. However, explicit models can provide initial points, serving as a stepping stone to make Gaussian splatting human mesh recovery from RGB inputs feasible. Animatable 3D Gaussian liu2023animatable learns human avatars from input images and poses by extending 3D Gaussian to dynamic human scenes. In the proposed framework, they model a set of skinned 3D Gaussian and a corresponding skeleton in canonical space and deform 3D Gaussian to posed space according to the input poses.
5.3 Summary of human mesh recovery
This chapter provides a comprehensive review of human mesh recovery utilizing both explicit and implicit models. Explicit models excel in robustly reconstructing human mesh but often fall short in capturing intricate details, prompting a range of extensions for greater detail accuracy. In contrast, implicit-based human mesh recovery tends to lack stability, which is known for its flexibility and adaptability. Consequently, several studies have explored integrating implicit models with explicit counterparts to synergize their respective strengths. The trade-off between flexibility and robustness in human mesh recovery represents a pivotal and enduring area of research.
6 Evaluation
6.1 Evaluation metrics
There are many evaluation metrics that can be used fairly to measure the performance of deep models in human pose estimation, and some of the key evaluation metrics are provided below.
Mean Per Joint Position Error (MPJPE) is widely used to evaluate the accuracy performance of 3D human pose estimation by calculating the L2 distance between the predicted joint coordinates and their ground truth counterparts. Denote the estimated coordinates of the j-th joint as $p_{j}^{*}$ and the ground truth coordinates as $p_{j}$. The MPJPE of $j$-th joint in the skeleton can be computed as:
$MPJPE=\frac{1}{N}\sum_{j=1}^{N}{\left\|p_{j}-p_{j}^{*}\right\|_{2}}$ (1)
where the skeleton comprises $N$ joints, and unlike previously in 2D, where the error is quantified in pixels, the joint coordinates in 3D are measured and reported in millimeters (mm).
Mean Per Joint Angle Error (MPJAE) quantifies the angular discrepancy between the estimated and groundtruth joints. The function $r_{j}^{*}$ returns the estimated angle of $j$-th joint, and function $r_{j}$ returns the ground truth angle. The MPJAE is computed in three dimensions as follows:
$MPJAE=\frac{1}{3N}\sum_{j=1}^{3N}{\left|\left(r_{j}-r_{j}^{*}\right)\mathrm{mod}\pm 180\right|}$ (2)
Mean Per Joint Localization Error (MPJLE) is a more perceptive and robust evaluation metric than MPJPE and MPJAE ionescu_human36m_2014 . It allows for an adjustable tolerance level through a perceptual threshold $t$ :
$MPJLE=\frac{1}{N}\sum_{j=1}^{N}{\mathds{1}_{\left\|l_{j}-l_{j}^{*}\right\|_{2}\geq t}}$ (3)
where $l(\cdot)$ indicates the joint localization.
There are various modifications of MPJPE, MPJAE, and MPJLE, including Procrustes-aligned (PA-) MPJPE, MPJAE, MPJLE and Normalized (N-) MPJPE, MPJAE, MPJLE. The former refers to procrustes-aligned metrics, while the latter denotes normalized metrics. Additionally, some 2D metrics are adaptable to 3D, such as the 3D Percentage of Correct Keypoints (3D PCK) and the 3D Area Under the Curve (3D AUC). The PCK YangandRamanan2013 measures the distance between predicted and groundtruth keypoints, considering keypoints correct if this distance is less than a predefined threshold. The AUC signifies the aggregate area beneath the PCK threshold curve as the threshold varies.
Mean Per Vertex Position Error (MPVPE) is a metric used to evaluate the human mesh reconstruction by computing the L2 distance between the predicted and ground truth mesh points. In some published articles, the MPVPE is also called Vertex-to-Vertex (V2V) error and Per-Vertex Error (PVE).
6.2 Datasets
Datasets are essential in developing deep learning-based human pose estimation and mesh recovery. Researchers have developed many datasets to train models and facilitate fair comparisons among different methods. In this section, we introduce the details of related datasets from recent years. Summary of these datasets, including information on 3D pose and 3D mesh, are presented in Table 3.
Table 3: The overview of the mainstream datasets.
Dataset Type Data Total frames Feature Download link
Human3.6M ionescu_human36m_2014 3D/Mesh Video 3.6M multi-view Website
3DPW von2018recovering 3D/Mesh Video 51K multi-person Website
MPI-INF-3DPH mehta2017monocular 2D/3D Video 2K in-wild Website
HumanEva sigal2010humaneva 3D Video 40K multi-view Website
CMU-Panoptic Joo_2015_ICCV 3D Video 1.5M multi-view/multi-person Website
MuCo-3DHP mehta2018single 3D Image 8K multi-person/occluded scence Website
SURREAL varol2017learning 2D/3D/Mesh Video 6.0M synthetic model Website
3DOH50K zhang2020object 2D/3D/Mesh Image 51K object-occluded Website
3DCP muller2021self Mesh Mesh 190 contact Website
AMASS mahmood2019amass Mesh Motion 11K soft-tissue dynamics Website
DensePose guler2018densepose Mesh Image 50K multi-person Website
UP-3D lassner2017unite 3D/Mesh Image 8K sport scence Website
THuman2.0 zheng2019deephuman Mesh Image 7K textured surface Website
Human3.6M dataset ionescu_human36m_2014 is the most widely used dataset in the evaluation of 3D human pose estimation. It encompasses a vast collection of 3.6 million poses captured using RGB and ToF cameras from diverse viewpoints within a real-world setting. Additionally, this dataset incorporates high-resolution 3D scanner data of body meshes, playing a pivotal role in the progression of human sensing systems. Table 4 exhibits the performance of state-of-the-art 3D human pose estimation methods on the Human3.6M dataset.
Table 4: Comparisons of 3D pose estimation methods on Human3.6M ionescu_human36m_2014 .
Method Year Publication Highlight MPJPEnormal-$\downarrow$ PMPJPEnormal-$\downarrow$ Code
Graformer zhao2022graformer 2022 CVPR’22 graph-based transformer 35.2 - Code
GLA-GCN yu2023gla 2023 ICCV’23 adaptive GCN 34.4 37.8 Code
PoseDA chai2023global 2023 arXiv’23 domain adaptation 49.4 34.2 Code
GFPose ci2023gfpose 2023 CVPR’23 gradient fields 35.6 30.5 Code
TP-LSTMs lee2022human 2022 TPAMI’22 pose similarity metric 40.5 31.8 -
FTCM tang2023ftcm 2023 TCSVT’23 frequency-temporal collaborative 28.1 - Code
VideoPose3D pavllo20193d 2019 CVPR’19 semi-supervised 46.8 36.5 Code
PoseFormer zheng20213d 2021 ICCV’21 spatio-temporal transformer 44.3 34.6 Code
STCFormer tang20233d 2023 CVPR’23 spatio-temporal transformer 40.5 31.8 Code
3Dpose_ssl wang20193d 2020 TPAMI’20 self-supervised 63.6 63.7 Code
MTF-Transformer shuai2022adaptive 2022 TPAMI’22 multi-view temporal fusion 26.2 - Code
AdaptPose gholami2022adaptpose 2022 CVPR’22 cross datasets 42.5 34.0 Code
3D-HPE-PAA xue2022boosting 2022 TIP’22 part aware attention 43.1 33.7 Code
DeciWatch zeng2022deciwatch 2022 ECCV’22 efficient framework 52.8 - Code
Diffpose gong2023diffpose 2023 CVPR’23 pose refine 36.9 28.7 Code
Elepose wandt2022elepose 2022 CVPR’22 unsupervised - 36.7 Code
Uplift and Upsample einfalt2023uplift 2023 CVPR’23 efficient transformers 48.1 37.6 Code
RS-Net hassan2023regular 2023 TIP’23 regular splitting graph network 48.6 38.9 Code
HSTFormer qian2023hstformer 2023 arXiv’23 spatial-temporal transformers 42.7 33.7 Code
PoseFormerV2 zhao2023poseformerv2 2023 CVPR’23 frequency domain 45.2 35.6 Code
DiffPose holmquist2023diffpose 2023 ICCV’23 diffusion models 42.9 30.8 Code
MPI-INF-3DPH dataset mehta2017monocular offers over 2K videos featuring joint annotations of 13 keypoints in outdoor scenes, apt for 2D and 3D human pose estimation. The ground truth was obtained using a multi-camera arrangement and a marker-less MoCap system, representing a shift from traditional marker-based MoCap systems that involve real individuals. Table 5 showcases the performance of state-of-the-art methods on the 3DPH dataset.
Table 5: Comparisons of 3D pose estimation methods on MPI-INF-3DPH mehta2017monocular .
Method Year Publication Highlight MPJPEnormal-$\downarrow$ PCKnormal-$\uparrow$ AUCnormal-$\uparrow$ Code
HSTFormer qian2023hstformer 2023 arXiv’23 spatial-temporal transformers 28.3 98.0 78.6 Code
PoseFormerV2 zhao2023poseformerv2 2023 CVPR’23 frequency domain 27.8 97.9 78.8 Code
Uplift and Upsample einfalt2023uplift 2023 CVPR’23 efficient transformers 46.9 95.4 67.6 Code
RS-Net hassan2023regular 2023 TIP’23 regular splitting graph network - 85.6 53.2 Code
Diffpose gong2023diffpose 2023 CVPR’23 pose refine 29.1 98.0 75.9 Code
FTCM tang2023ftcm 2023 TCSVT’23 frequency-temporal collaborative 31.2 97.9 79.8 Code
STCFormer tang20233d 2023 CVPR’23 spatio-temporal transformer 23.1 98.7 83.9 Code
PoseDA chai2023global 2023 arXiv’23 domain adaptation 61.3 92.0 62.5 Code
TP-LSTMs lee2022human 2022 TPAMI’22 pose similarity metric 48.8 82.6 81.3 -
AdaptPose gholami2022adaptpose 2022 CVPR’22 cross datasets 77.2 88.4 54.2 Code
3D-HPE-PAA xue2022boosting 2022 TIP’22 part aware attention 69.4 90.3 57.8 Code
Elepose wandt2022elepose 2022 CVPR’22 unsupervised 54.0 86.0 50.1 Code
3DPW dataset von2018recovering captures 51,000 sequences of single-view video, complemented by IMUs data. These videos were recorded using a handheld camera, with the IMU data facilitating the association of 2D poses with their 3D counterparts. 3DPW stands out as one of the most formidable datasets, establishing itself as a benchmark for 3D pose estimation in multi-person, in-wild scenarios of recent times. Table 6 shows the performance of state-of-the-art human mesh recovery methods on the Human3.6M and 3DPW datasets.
Table 6: Comparisons of human mesh recovery methods on Human3.6M ionescu_human36m_2014 and 3DPW von2018recovering .
Method Publication Highlight Human3.6M 3DPW Code
MPJPEnormal-$\downarrow$ PA-MPJPEnormal-$\downarrow$ MPJPEnormal-$\downarrow$ PA-MPJPEnormal-$\downarrow$ PVEnormal-$\downarrow$
VirtualMarker ma20233d CVPR’23 novel intermediate representation 47.3 32.0 67.5 41.3 77.9 Code
NIKI li2023niki CVPR’23 inverse kinematics - - 71.3 40.6 86.6 Code
TORE dou2023tore ICCV’23 efficient transformer 59.6 36.4 72.3 44.4 88.2 Code
JOTR li2023jotr ICCV’23 contrastive learning - - 76.4 48.7 92.6 Code
HMDiff foo2023distribution ICCV’23 reverse diffusion processing 49.3 32.4 72.7 44.5 82.4 Code
ReFit wang2023refit ICCV’23 recurrent fitting network 48.4 32.2 65.8 41.0 - Code
PyMAF-X zhang2023pymaf TPAMI’23 regression-based one-stage whole body - - 74.2 45.3 87.0 Code
PointHMR kim2023sampling CVPR’23 vertex-relevant feature extraction 48.3 32.9 73.9 44.9 85.5 -
PLIKS shetty2023pliks CVPR’23 inverse kinematics 47.0 34.5 60.5 38.5 73.3 Code
ProPose fang2023learning CVPR’23 learning analytical posterior probability 45.7 29.1 68.3 40.6 79.4 Code
POTTER zheng2023potter CVPR’23 pooling attention transformer 56.5 35.1 75.0 44.8 87.4 Code
PoseExaminer liu2023poseexaminer ICCV’23 automated testing of out-of-distribution - - 74.5 46.5 88.6 Code
MotionBERT zhu2023motionbert ICCV’23 pretrained human representations 43.1 27.8 68.8 40.6 79.4 Code
3DNBF zhang20233d ICCV’23 analysis-by-synthesis approach - - 88.8 53.3 - Code
FastMETRO cho2022cross ECCV’22 efficient architecture 52.2 33.7 73.5 44.6 84.1 Code
CLIFF li2022cliff ECCV’22 multi-modality inputs 47.1 32.7 69.0 43.0 81.2 Code
PARE kocabas2021pare ICCV’21 part-driven attention - - 74.5 46.5 88.6 Code
Graphormer lin2021mesh ICCV’21 GCNN-reinforced transformer 51.2 34.5 74.7 45.6 87.7 Code
PSVT qiu2023psvt CVPR’23 spatio-temporal encoder - - 73.1 43.5 84.0 -
GLoT shen2023global CVPR’23 short-term and long-term temporal correlations 67.0 46.3 80.7 50.6 96.3 Code
MPS-Net wei2022capturing CVPR’23 temporally adjacent representations 69.4 47.4 91.6 54.0 109.6 Code
MAED wan2021encoder ICCV’21 multi-level attention 56.4 38.7 79.1 45.7 92.6 Code
Lee et al. lee2021uncertainty ICCV’21 uncertainty-aware 58.4 38.4 92.8 52.2 106.1 -
TCMR choi2021beyond CVPR’21 temporal consistency 62.3 41.1 95.0 55.8 111.3 -
VIBE kocabas2020vibe CVPR’20 self-attention temporal network 65.6 41.4 82.9 51.9 99.1 Code
ImpHMR cho2023implicit CVPR’23 implicitly imagine person in 3D space - - 74.3 45.4 87.1 -
SGRE wang20233d ICCV’23 sequentially global rotation estimation - - 78.4 49.6 93.3 Code
PMCE you2023co ICCV’23 pose and mesh co-evolution network 53.5 37.7 69.5 46.7 84.8 Code
HumanEva dataset sigal2010humaneva is a multi-view 3D human pose estimation dataset comprising two versions: HumanEva-I and HumanEva-II. In HumanEva-I, the dataset includes around 40,000 multi-view video frames captured from seven cameras positioned at the front, left, and right (RGB) and four corners (Mono). Additionally, HumanEva-II features approximately 2,460 frames, recorded with four cameras at each corner.
CMU-Panoptic dataset Joo_2015_ICCV ; simon2017hand includes 65 frame sequences, approximately 5.5 hours of footage, and features 1.5 million 3D annotated poses. Recorded via a massively multi-view system equipped with 511 calibrated cameras and 10 RGB-D sensors featuring hardware-based synchronization, this dataset is crucial for developing weakly supervised methods through multi-view geometry. These methods address the occlusion problems commonly encountered in traditional computer vision techniques.
Multiperson Composited 3D Human Pose (MuCo-3DHP) dataset mehta2018single serves as a large-scale, multi-person occluded training set for 3D human pose estimation. Frames in the MuCo-3DHP are generated from the MPI-INF-3DPH dataset through a compositing and augmentation scheme.
SURREAL dataset varol2017learning is a large synthetic human body dataset containing 6 million RGB video frames. It provides a range of accurate annotations, including depth, body parts, optical flow, 2D/3D poses, and surfaces. In the SURREAL dataset, images exhibit variations in texture, view, and pose, and the body models are based on the SMPL parameters, a widely-recognized mesh representation standard.
3DOH50K dataset zhang2020object offers a collection of 51,600 images obtained from six distinct viewpoints in real-world settings, predominantly featuring object occlusions. Each image is annotated with ground truth 2D and 3D poses, SMPL parameters, and a segmentation mask. Utilized for training human estimation and reconstruction models, the 3DOH50K dataset facilitates exceptional performance in occlusion scenarios.
3DCP dataset muller2021self represents a 3D human mesh dataset, derived from AMASS mahmood2019amass . It includes 190 self-contact meshes spanning six human subjects (three males and three females), each modeled with an SMPL-X parameterized template.
AMASS dataset mahmood2019amass constitutes a comprehensive and diverse human motion dataset, encompassing over 11,000 motions from 300 subjects, totaling more than 40 hours. The motion data, accompanied by SMPL parameters for skeleton and mesh representation, is derived from a marker-based MoCap system utilizing 15 optical markers.
DensePose dataset guler2018densepose features 50,000 manually annotated real images, comprising 5 million image-to-surface correspondence pairs extracted from the COCO lin2014microsoft dataset. This dataset proves instrumental for training in dense human pose estimation, as well as in detection and segmentation tasks.
UP-3D dataset lassner2017unite is a dedicated 3D human pose and shape estimation dataset featuring extensive annotations in sports scenarios. The UP-3D comprises approximately 8,000 images from the LSP and MPII datasets. Additionally, each image in UP-3D is accompanied by a metadata file indicating the quality (medium or high) of the 3D fit.
THuman dataset zheng2019deephuman constitutes a 3D real-world human mesh dataset. It includes 7,000 RGBD images, each featuring a textured surface mesh obtained using a Kinect camera. Including surface mesh with detailed texture and the aligned SMPL model is anticipated to significantly enhance and stimulate future research in human mesh reconstruction.
7 Applications
In this section, we review related works about human pose estimation and mesh recovery for a few popular applications.
Motion Retargeting: Human motion retargeting can transfer human actions onto actors. Using pose estimation eliminates the need for motion capture systems and achieves image-to-image translation. Therefore, 3D human pose estimation is crucial for retargeting. Recently, there has been a significant amount of retargeting work based on 3D human pose estimation aberman2020skeleton ; yang2020transmomo ; yu2023bidirectionally . End-to-end methods and related datasets are also designed gomes2021shape . Additionally, unsupervised methods in In-the-Wild scenarios zhu2022mocanet have been developed, achieved through canonicalization operations and derived regularizations. Beyond bodily retargeting, facial retargeting mo2022towards ; chen2024morphable has also gained prominence, wherein the intricacies of facial expressions offer a refined portrayal of the actor’s emotional and psychological states.
Action Recognition: Action recognition employs algorithms to identify and analyze human movements from images or videos. The results derived from 3D human pose estimation and reconstruction play a pivotal role in deciphering the dynamics of human motion within a three-dimensional context, thereby transforming these movements into actionable behavioral insights yang2021feedback ; mazzia2022action ; duan2022revisiting ; lu2023hard . Moreover, these advancements can significantly augment the efficiency of action recognition bian2021structural . It is also possible to address pose estimation and action recognition within the same framework through multi-task learning and sharing features luvizon2020multi .
Security Monitoring: In video surveillance systems at public places or critical facilities, pedestrian tracking and re-identification are key tasks. Combined with human pose estimation, Human tracking is utilized for the surveillance and analysis of pedestrian flow, tracking specific targets, and tracking and analyzing human behavior in space. This approach is highly beneficial for tracking humans in complex scenarios bao2020pose ; reddy2021tessetrack ; goel2023humans ; sun2023trace . Considering the constrained viewpoints of individual cameras in such systems, Re-identification enables the recognition and tracking of the same person as they move across different camera views. Human pose estimation significantly contributes to the enhancement of this re-identification wang2021horeid .
SLAM: SLAM precisely estimates its location by gathering sensor data from its environment, employing technologies such as cameras and LiDAR. It simultaneously constructs or refines an environmental map, facilitating self-localization and map creation in unfamiliar territories. Diverging from the conventional SLAM systems that predominantly concentrate on objects, recent advancements have shifted focus towards incorporating humans within the environmental context. Notably, Dai et al. dai2021indoor have illustrated the significance of 3D human trajectory reconstruction in indoor settings, enhancing indoor navigation capabilities. Furthermore, Kocabas et al. kocabas2023pace have innovatively integrated human motion priors into SLAM, effectively merging human pose estimation with scene analysis.
Autonomous Driving: In robotic navigation and autonomous vehicle applications, estimating human pose enables these systems to comprehend human behavior and intentions better. This understanding facilitates more intelligent decision-making and interaction. Zheng et al. zheng2022multi propose a multi-modal approach employing 2D labels on RGB images as weak supervision for 3D human pose estimation in the context of autonomous vehicles. Wang et al. wang2023learning have developed a comprehensive framework to learn physically plausible human dynamics from real driving scenarios, effectively bridging the gap between actual and simulated human behavior in safety-critical applications.
Human–Computer Interaction: Human-computer interaction entails the bidirectional communication between humans and computers, underscored by the critical need for computers to interpret human poses accurately. Liu et al. liu2022arhpe advanced this field by proposing asymmetric relation-aware representation learning for head pose estimation in industrial human–computer interaction, which utilizes an effective Lorentz distribution learning scheme. In Augmented Reality (AR) systems, the accuracy of human pose estimation significantly elevates the interaction quality between virtual objects and the real environment, fostering more natural interactions between humans and virtual entities. In this context, Weng et al. weng2019photo developed a novel method and application for animating a human subject from a single photo within AR.
8 Challenges and Conclusion
In this survey, we have presented a contemporary overview of recent deep learning-based 3D human pose estimation and mesh recovery methods. A comprehensive taxonomy and performance comparison of these methods has been covered. We further point out a few promising research directions, hoping to promote advances in this field.
Large Models. The efficacy of large models in language anil2023palm ; achiam2023gpt and foundational computer vision tasks, such as segmentation kirillov2023segment and tracking yang2023track , has been universally acknowledged and is quite remarkable. Additionally, while there is burgeoning research in human-centric computer vision tasks ci2023unihcp , the area of 3D tasks still demands further investigation. Furthermore, the development of large models not only constitutes a significant area of research but also the exploration of their effective utilization presents substantial challenges and practical importance. Exploratory efforts to fuse pose estimation with large language models feng2023posegpt and the combination of large visual models with pose estimation are also meaningful.
More Detailed Reconstruction. At present, explicit model-based methodologies such as SMPL loper2015smpl and SMPL-X cai2023smpler , fall short of meeting the demands for detail that people expect. On the other hand, implicit representation approaches saito2019pifu ; huang2020arch ; he2020geo , as well as rendering techniques such as NeRF gao2022mps ; mu2023actorsnerf and Gaussian Splatting liu2023animatable , are capable of capturing fine details but lack sufficient robustness in pose estimation. Bridging the gap between robust pose estimation and surface details remains a formidable challenge that necessitates collaborative efforts from computer vision and computer graphics researchers.
Crowding and Occlusion Challenges. In open-world scenarios, the phenomena of crowding and occlusion are prevalent, and they represent long-standing challenges in the field of object detection. Currently, top-down methods depend on object detection, rendering these issues inescapable. Although bottom-up strategies may circumvent object detection, they confront formidable challenges regarding key point assembly.
Speed. Speed is an essential aspect to consider in the practical deployment of algorithms. While most current research papers report achieving real-time performance on GPUs, a wide array of applications necessitates real-time and efficient processing on edge computing platforms, notably on ARM processors within smartphones. The disparity in performance between ARM processors and GPUs is pronounced, underscoring the immense value of optimizing for speed.